# Geometric integrators for adiabatically closed simple thermodynamic systems

### v2

@@ 1 Introduction
The use of differential geometric methods in the study of thermodynamics dates back considerably. Traditionally, the geometry of equilibrium thermodynamics has been mainly studied via contact geometry. In this setting, thermodynamic properties are encoded by Legendre submanifolds of the thermodynamic phase space (see [bravetti0, bravetti2019, lacomba1, lacomba2, mrugala1, mrugala2] and references therein). In [A.d.L+2020], a similar approach was considered, but using the so-called evolution vector field instead of the contact Hamiltonian vector field.
@@
Recently, Gay-Balmaz and Yoshimura [G.Y2018, Y.G2022] introduced variational principles for the description of thermodynamic systems. Their formulation extends the Hamilton principle of classical mechanics to include irreversible processes by introducing additional phenomenological and variational constraints. In a recent paper [d.B2025], the first two authors of the present article introduced almost cosymplectic structures, new geometric structures that lead to the same dynamics as Gay-Balmaz and Yoshimura’s approach. An almost cosymplectic structure on a $(2n+1)$-dimensional manifold $M$ is a pair $(\omega,\eta)$ consisting of a $2$-form $\omega$ and a $1$-form $\eta$ such that $\omega^{n}\wedge\eta$ is a volume form on $M$. These structures were already considered in [BracketsDuales], albeit in a different context. They are just a slight generalisation of cosymplectic structures, which were introduced by Libermann [libermann]. A cosymplectic structure is simply an almost cosymplectic structure $(\omega,\eta)$ such that both $\omega$ and $\eta$ are closed forms. We shall be considering the so-called partially cosymplectic structures, for which $\omega$ is assumed to be closed, but $\eta$ is not. Note that any contact form $\eta$ defines a partially cosymplectic structure $(\mathrm{d}\eta,\eta)$.
@@
In the current paper, we take advantage of this geometric description to obtain a discrete description and construct geometric integrators appropriate for treating these thermodynamic systems. Due to the complexity of the different types of systems, we focus on adiabatically closed simple thermodynamic systems. We hope to develop similar procedures for more complex systems, and thus cover the entire spectrum of cases, in a forthcoming publication. Our discrete model is inspired by variational integrators (see the quintessential review by Marsden and West [M.W2001]), as well as their extensions to contact Lagrangian systems [A.M.L+2021, V.B.S2019].
@@
In [G.Y2018a], Gay-Balmaz and Yoshimura also developed variational integrators for the nonequilibrium thermodynamics of simple closed systems, following a similar approach to ours. Let us not forget that this type of discretization of systems described by a Lagrangian has been studied for more than five decades (see [M.W2001] and references therein). The main difference between our development and that of Gay-Balmaz and Yoshimura is not only that they consider two entropies (there are no major differences in this regard), but also that our approach is based on the geometric description of the thermodynamic system, which allows us to use the standard techniques of geometric mechanics. Here we have addressed only the case of infinitesimal symmetries and Noether’s theorem, as well as the existence of a momentum mapping, but in subsequent papers we propose to address aspects such as coisotropic reduction, Hamilton–Jacobi theory, and the extension to more complex systems. These objectives require a thorough understanding of the underlying geometry, and that is the first step we are currently developing. A more explicit comparison between our approach and that of Gay-Balmaz and Yoshimura can be found in Table 1.
@@
It is also worth mentioning the numerical methods for a supercritical van der Waals fluid enclosed in a cylinder developed by Fülöp, Szücz and Takács in [F.S.T2025]. Their approach is based on the GENERIC framework, which essentially consists of expressing a dissipative dynamical system in the form
@@
where $(L^{i}_{j})$ is a skew-symmetric matrix (representing a presymplectic form), $(M^{i}_{j})$ is a symmetric matrix, while $E$ and $S$ are functions that can be regarded as the total energy and the total entropy of the system, respectively. The authors then consider a generalisation of the symplectic Euler scheme, to discretise the evolution equations of the fluid. They obtain meritorious numerical results. However, they do not develop a general method for discretising thermodynamic systems, but rather construct an ad hoc generalisation of the symplectic method that serves their system under consideration. Moreover, their approach does not consider the geometry of the thermodynamic phase space whatsoever. We would also like to mention that the dynamics (1) corresponds to a Rayleigh-type dissipation, i.e. a dissipation proportional to velocity or momentum. Indeed, the symmetric matrix $(M^{i}_{j})$ determines a Rayleigh tensor. We previously studied systems with Rayleigh dissipation in [d.L.L2021, d.L.L2022] and their discretisation in [d.L.L2022a].
@@ 2 Preliminaries
In this section, we will present the notion of partially cosymplectic structures introduced in [BracketsDuales, d.B2025], and recall the continuous model for an adiabatically closed simple thermodynamic system (see [G.Y2018, d.B2025]).
@@ 2.1 Partially cosymplectic structures
Notice that a top-degree form on a manifold is a volume form if and only if its kernel is the zero-section. The equivalence between the first and second statements then follows from the fact that
$\ker\left(\omega^{n}\wedge\eta\right)=\ker\omega\cap\ker\eta\,.$
@@
A $1$-form $\eta$ on a $(2n+1)$-dimensional manifold $M$ is a contact form if and only if the pair $(\mathrm{d}\eta,\eta)$ is a partially cosymplectic structure on $M$. The Reeb vector field with respect to both structures is the same one; and the isomorphism between ${\mathsf{T}}M$ and ${\mathsf{T}}^{\ast}M$ defined by $\eta$ is precisely $\flat_{(\eta,\,\mathrm{d}\eta)}$. For more details, see [de2019contact]. ∎
@@
For additional details on almost and partially cosymplectic structures, refer to [BracketsDuales, d.B2025].
@@ 2.2 The continuous model: adiabatically closed simple thermodynamic systems
The phase space of a mechanical Hamiltonian system can usually be regarded as the cotangent bundle ${\mathsf{T}}^{\ast}Q$ of its configuration space $Q$. This means that the position of the system is determined by a point $q\in Q$, and its momentum is a covector on the corresponding fiber: $p_{q}\in{\mathsf{T}}^{\ast}_{q}Q$. As it is well-known, ${\mathsf{T}}^{\ast}Q$ is canonically endowed with a symplectic form $\omega_{Q}$. Indeed, $\omega_{\mathcal{Q}}=-d\Theta_{Q}$, where $\Theta_{Q}=p_{i}\,dq^{i}$ is the Liouville form on ${\mathsf{T}}^{\ast}{Q}$ in bundle coordinates $(q^{i},p_{i})$ on ${\mathsf{T}}^{\ast}Q$ associated with some local coordinates $(q^{i})$ on $Q$. Then, we have $\omega_{Q}=\mathrm{d}q^{i}\wedge\mathrm{d}p_{i}$ (see [godbillon1969geometrie, abraham, de2011methods]).
@@
Assuming that ${\frac{\partial H}{\partial S}}$ is nowhere-vanishing, then the pair $(\omega,\eta)$ is a partially cosymplectic structure on $M$, regardless of what expression $F^{\mathrm{fr}}$ takes. As we shall explain below, for the type of Hamiltonians that we will consider, ${\frac{\partial H}{\partial S}}$ will be the temperature of the system. Hence, our assumption simply means that the temperature of the system is not the absolute zero. It is worth recalling that by the third law of thermodynamics, the absolute zero cannot be reached adiabatically [Kieu2019]. Hereinafter, we shall thus assume that $(M,\omega,\eta)$ is a partially cosymplectic manifold. The tuple $(M,\omega,\eta,H,F^{\mathrm{ext}},F^{\mathrm{fr}})$ will be called the Hamiltonian thermodynamic system.
@@
It is noteworthy that equations (11) are equivalent to those obtained by Gay-Balmaz and Yoshimura [G.Y2018, Y.G2022].
@@
Following [G.Y2018], in this formulation, the temperature is defined as minus the derivative of $L$ with respect to $S$, i.e., $T=-\frac{\partial L}{\partial S}$, which is assumed to be positive. When the Lagrangian has the standard form
@@
Previous studies of thermodynamics have been carried out using contact geometry. In [A.d.L+2020], given a function $H\in\mathscr{C}^{\infty}(M)$, a contact evolution field $Y_{H}\in\mathfrak{X}(M)$ is defined as
@@
On the other hand, we can endow $M$ with the partially cosymplectic structure $(\omega,\eta)$ given by (5), for $F^{\mathrm{fr}}_{i}=-\mathcal{R}_{C}(H)p_{i}$. In this way, we can write $Y_{H}=\mathcal{E}_{H,F^{\mathrm{ext}}}$, with $F^{\mathrm{ext}}\equiv 0$. This can be easily verified by comparing the local expression of $Y_{H}$ above with (7), the local expression of $\mathcal{E}_{H,F^{\mathrm{ext}}}$. Thus, the evolution vector defined in [A.d.L+2020] can be regarded as a particular case of the one defined in this paper. Notice that, since $\eta\neq\eta_{C}$, the isomorphisms $\flat_{(\eta_{C},\,\mathrm{d}\eta_{C})}$ and $\flat_{(\eta,\,\mathrm{d}\eta)}$ are different.
Similarly, the thermodynamical Herglotz equations for a contact Lagrangian from [A.d.L+2020] can be recovered from our thermodynamic Euler–Lagrange equations (11). Indeed, taking $\tilde{F}^{\mathrm{fr}}_{i}=\frac{\partial L}{\partial S}\frac{\partial L}{\partial v_{i}}$ and $\tilde{F}^{\mathrm{ext}}\equiv 0$, we have that
@@
is the so-called stable tangent structure on $\mathbb{R}\times{\mathsf{T}}Q$ (see [poincare]). Such a vector field is called a Non-autonomous Second Order Differential Equations (NSODE, for short) and its local expression is as follows:
@@
where $L\colon{\mathsf{T}}Q\times\mathbb{R}\to\mathbb{R}$ is the Lagrangian function (see [A.d.L+2020]). However, in the current paper that function $\mathcal{A}$ is more general since the system can also include external and friction forces.
@@
See [yano-ishihara, de2011methods] for the intrinsic definitions and additional details.
@@ 3.1 Discrete variational principle
Notice that the equations defining the discrete thermodynamic path space are a discretization of the phenomenological constraints used in [G.Y2018]. We should also observe that, in general, they depend on the coordinate chart chosen, since $f^{\mathrm{fr,\pm}}_{d}q_{k}$ does. However, if $Q$ is a vector space, we can choose linear coordinates and take $f^{\mathrm{fr,+}}_{d}=f^{\mathrm{fr,-}}_{d}$. In that case, the definition above becomes coordinate-independent. ∎
@@
Equation (21) can be interpreted as the discretization of the variational constraint introduced in [G.Y2018], via the identification
@@ 3.2 Discrete flows
These discrete thermodynamic Legendre transforms allows us to characterise the solutions of the discrete thermodynamic Euler–Lagrange equations via momentum matching equations as in [M.W2001]. Indeed, definition 7 and equations (22) immediatly imply the following:
@@ 3.3 Discrete Symmetries and Discrete Noether’s theorem
The proof, mutatis mutandis, is identical to the one from [d.L.L2022, Theorem 2].
@@
Table 1: Comparison between the discretisation provided by our method and the one developed in [G.Y2018a] for a Lagrangian thermodynamic system $({\mathsf{T}}Q\times\mathbb{R},\omega_{L},\eta_{L},L,\tilde{F}^{\mathrm{ext}},\tilde{F}^{\mathrm{fr}})$. Note that in their approach the evolution is given by an equation and a certain constraint $C_{d}\subseteq Q\times Q\times\mathbb{R}\times\mathbb{R}$.
@@ 4.1 Damped harmonic oscillator
Let us consider a one dimensional damped harmonic oscillator, such as in [A.d.L+2020], whose continuous Lagrangian $L:{\mathsf{T}}\mathbb{R}\times\mathbb{R}\cong\mathbb{R}^{3}\to\mathbb{R}$ is given by
@@ 4.2 Ideal gas in a piston
Consider a perfect gas confined by a piston and contained in a cylinder, as was done by Gay-Balmaz and Yoshimura in [Balmaz+2017]. In suitable units and for a suitable number of particles in the gas, the Lagrangian that describes this system is given by
@@ 4.3 Van der Waals gas in a piston
A more realistic example would be to consider a Van der Waals gas formed by particles without internal degrees of freedom, contained in a cylinder similar to that in the previous example. Considering a suitable value for the mass of the cylinder, as well as for the number of particles in the gas, the Lagrangian in suitable units is given by [Johnston2014]:

### v3

@@ 1 Introduction
The use of differential geometric methods in the study of thermodynamics dates back considerably. Traditionally, the geometry of equilibrium thermodynamics has been mainly studied via contact geometry. In this setting, thermodynamic properties are encoded by Legendre submanifolds of the thermodynamic phase space (see [5, 6, 20, 21, 28, 29] and references therein). In [2], a similar approach was considered, but using the so-called evolution vector field instead of the contact Hamiltonian vector field.
@@
Recently, Gay-Balmaz and Yoshimura [16, 33] introduced variational principles for the description of thermodynamic systems. Their formulation extends the Hamilton principle of classical mechanics to include irreversible processes by introducing additional phenomenological and variational constraints. In a recent paper [10], the first two authors of the present article introduced almost cosymplectic structures, new geometric structures that lead to the same dynamics as Gay-Balmaz and Yoshimura’s approach. An almost cosymplectic structure on a $(2n+1)$-dimensional manifold $M$ is a pair $(\omega,\eta)$ consisting of a $2$-form $\omega$ and a $1$-form $\eta$ such that $\omega^{n}\wedge\eta$ is a volume form on $M$. These structures were already considered in [22], albeit in a different context. They are just a slight generalisation of cosymplectic structures, which were introduced by Libermann [25]. A cosymplectic structure is simply an almost cosymplectic structure $(\omega,\eta)$ such that both $\omega$ and $\eta$ are closed forms. We shall be considering the so-called partially cosymplectic structures, for which $\omega$ is assumed to be closed, but $\eta$ is not. Note that any contact form $\eta$ defines a partially cosymplectic structure $(\mathrm{d}\eta,\eta)$.
@@
In the current paper, we take advantage of this geometric description to obtain a discrete description and construct geometric integrators appropriate for treating these thermodynamic systems. Due to the complexity of the different types of systems, we focus on adiabatically closed simple thermodynamic systems. We hope to develop similar procedures for more complex systems, and thus cover the entire spectrum of cases, in a forthcoming publication. Our discrete model is inspired by variational integrators (see the quintessential review by Marsden and West [27]), as well as their extensions to contact Lagrangian systems [3, 31].
@@
In [17], Gay-Balmaz and Yoshimura also developed variational integrators for the nonequilibrium thermodynamics of simple closed systems, following a similar approach to ours. Let us not forget that this type of discretization of systems described by a Lagrangian has been studied for more than five decades (see [27] and references therein). The main difference between our development and that of Gay-Balmaz and Yoshimura is not only that they consider two entropies (there are no major differences in this regard), but also that our approach is based on the geometric description of the thermodynamic system, which allows us to use the standard techniques of geometric mechanics. Here we have addressed only the case of infinitesimal symmetries and Noether’s theorem, as well as the existence of a momentum mapping, but in subsequent papers we propose to address aspects such as coisotropic reduction, Hamilton–Jacobi theory, and the extension to more complex systems. These objectives require a thorough understanding of the underlying geometry, and that is the first step we are currently developing. A more explicit comparison between our approach and that of Gay-Balmaz and Yoshimura can be found in Table 1.
@@
It is also worth mentioning the numerical methods for a supercritical van der Waals fluid enclosed in a cylinder developed by Fülöp, Szücz and Takács in [14]. Their approach is based on the GENERIC framework, which, as its name suggests, is a very general framework for describing systems with dissipation (see [30]). More precisely, the authors consider a dissipative dynamical system of the form
@@
where $(L^{i}_{j})$ is a skew-symmetric matrix (representing a presymplectic form), $(M^{i}_{j})$ is a symmetric matrix, while $E$ and $S$ are functions that can be regarded as the total energy and the total entropy of the system, respectively. The authors then consider a generalisation of the symplectic Euler scheme, to discretise the evolution equations of the fluid. They obtain meritorious numerical results. However, they do not develop a general method for discretising thermodynamic systems, but rather construct an ad hoc generalisation of the symplectic method that serves their system under consideration. Moreover, their approach does not consider the geometry of the thermodynamic phase space whatsoever. We would also like to mention that the dynamics (1) corresponds to a Rayleigh-type dissipation, i.e. a dissipation proportional to velocity or momentum. Indeed, the symmetric matrix $(M^{i}_{j})$ determines a Rayleigh tensor. We previously studied systems with Rayleigh dissipation in [11, 13] and their discretisation in [12].
@@ 2 Preliminaries
In this section, we will present the notion of partially cosymplectic structures introduced in [22, 10], and recall the continuous model for an adiabatically closed simple thermodynamic system (see [16, 10]).
@@ 2.1 Partially cosymplectic structures
Notice that a top-degree form on a manifold is a volume form if and only if its kernel is the zero-section. The equivalence between the first and second statements then follows from the equivalence between the following statements:
  1. i)
    $\iota_{v}\left(\omega^{n}\wedge\eta\right)=0\Rightarrow v=0$,
  2. ii)
    $\iota_{v}\omega=0\text{ and }\iota_{v}\eta=0\Rightarrow v=0$.
@@
A $1$-form $\eta$ on a $(2n+1)$-dimensional manifold $M$ is a contact form if and only if the pair $(\mathrm{d}\eta,\eta)$ is a partially cosymplectic structure on $M$. The Reeb vector field with respect to both structures is the same one; and the isomorphism between ${\mathsf{T}}M$ and ${\mathsf{T}}^{\ast}M$ defined by $\eta$ is precisely $\flat_{(\eta,\,\mathrm{d}\eta)}$. For more details, see [8]. ∎
@@
For additional details on almost and partially cosymplectic structures, refer to [22, 10].
@@ 2.2 The continuous model: adiabatically closed simple thermodynamic systems
The phase space of a mechanical Hamiltonian system can usually be regarded as the cotangent bundle ${\mathsf{T}}^{\ast}Q$ of its configuration space $Q$. This means that the position of the system is determined by a point $q\in Q$, and its momentum is a covector on the corresponding fiber: $p_{q}\in{\mathsf{T}}^{\ast}_{q}Q$. As it is well-known, ${\mathsf{T}}^{\ast}Q$ is canonically endowed with a symplectic form $\omega_{Q}$. Indeed, $\omega_{\mathcal{Q}}=-d\Theta_{Q}$, where $\Theta_{Q}=p_{i}\,dq^{i}$ is the Liouville form on ${\mathsf{T}}^{\ast}{Q}$ in bundle coordinates $(q^{i},p_{i})$ on ${\mathsf{T}}^{\ast}Q$ associated with some local coordinates $(q^{i})$ on $Q$. Then, we have $\omega_{Q}=\mathrm{d}q^{i}\wedge\mathrm{d}p_{i}$ (see [19, 1, 9]).
@@
Assuming that ${\frac{\partial H}{\partial S}}$ is nowhere-vanishing, then the pair $(\omega,\eta)$ is a partially cosymplectic structure on $M$, regardless of what expression $F^{\mathrm{fr}}$ takes. As we shall explain below, for the type of Hamiltonians that we will consider, ${\frac{\partial H}{\partial S}}$ will be the temperature of the system. Hence, our assumption simply means that the temperature of the system is not the absolute zero. It is worth recalling that by the third law of thermodynamics, the absolute zero cannot be reached adiabatically [24]. Hereinafter, we shall thus assume that $(M,\omega,\eta)$ is a partially cosymplectic manifold. The tuple $(M,\omega,\eta,H,F^{\mathrm{ext}},F^{\mathrm{fr}})$ will be called the Hamiltonian thermodynamic system.
@@
It is noteworthy that equations (11) are equivalent to those obtained by Gay-Balmaz and Yoshimura [16, 33].
@@
Following [16], in this formulation, the temperature is defined as minus the derivative of $L$ with respect to $S$, i.e., $T=-\frac{\partial L}{\partial S}$, which is assumed to be positive. When the Lagrangian has the standard form
@@
Previous studies of thermodynamics have been carried out using contact geometry. In [2], given a function $H\in\mathscr{C}^{\infty}(M)$, a contact evolution field $Y_{H}\in\mathfrak{X}(M)$ is defined as
@@
On the other hand, we can endow $M$ with the partially cosymplectic structure $(\omega,\eta)$ given by (5), for $F^{\mathrm{fr}}_{i}=-\mathcal{R}_{C}(H)p_{i}$. In this way, we can write $Y_{H}=\mathcal{E}_{H,F^{\mathrm{ext}}}$, with $F^{\mathrm{ext}}\equiv 0$. This can be easily verified by comparing the local expression of $Y_{H}$ above with (7), the local expression of $\mathcal{E}_{H,F^{\mathrm{ext}}}$. Thus, the evolution vector defined in [2] can be regarded as a particular case of the one defined in this paper. Notice that, since $\eta\neq\eta_{C}$, the isomorphisms $\flat_{(\eta_{C},\,\mathrm{d}\eta_{C})}$ and $\flat_{(\eta,\,\mathrm{d}\eta)}$ are different.
Similarly, the thermodynamical Herglotz equations for a contact Lagrangian from [2] can be recovered from our thermodynamic Euler–Lagrange equations (11). Indeed, taking $\tilde{F}^{\mathrm{fr}}_{i}=\frac{\partial L}{\partial S}\frac{\partial L}{\partial v_{i}}$ and $\tilde{F}^{\mathrm{ext}}\equiv 0$, we have that
@@
is the so-called stable tangent structure on $\mathbb{R}\times{\mathsf{T}}Q$ (see [7]). Such a vector field is called a Non-autonomous Second Order Differential Equations (NSODE, for short) and its local expression is as follows:
@@
where $L\colon{\mathsf{T}}Q\times\mathbb{R}\to\mathbb{R}$ is the Lagrangian function (see [2]). However, in the current paper that function $\mathcal{A}$ is more general since the system can also include external and friction forces.
@@
See [32, 9] for the intrinsic definitions and additional details.
@@ 3.1 Discrete variational principle
Notice that the equations defining the discrete thermodynamic path space are a discretization of the phenomenological constraints used in [16]. We should also observe that, in general, they depend on the coordinate chart chosen, since $f^{\mathrm{fr,\pm}}_{d}q_{k}$ does. However, if $Q$ is a vector space, we can choose linear coordinates and take $f^{\mathrm{fr,+}}_{d}=f^{\mathrm{fr,-}}_{d}$. In that case, the definition above becomes coordinate-independent. ∎
@@
Equation (21) can be interpreted as the discretization of the variational constraint introduced in [16], via the identification
@@ 3.2 Discrete flows
These discrete thermodynamic Legendre transforms allows us to characterise the solutions of the discrete thermodynamic Euler–Lagrange equations via momentum matching equations as in [27]. Indeed, definition 7 and equations (22) immediatly imply the following:
@@ 3.3 Discrete Symmetries and Discrete Noether’s theorem
The proof, mutatis mutandis, is identical to the one from [12, Theorem 2].
@@
Table 1: Comparison between the discretisation provided by our method and the one developed in [17] for a Lagrangian thermodynamic system $({\mathsf{T}}Q\times\mathbb{R},\omega_{L},\eta_{L},L,\tilde{F}^{\mathrm{ext}},\tilde{F}^{\mathrm{fr}})$. Note that in their approach the evolution is given by an equation and a certain constraint $C_{d}\subseteq Q\times Q\times\mathbb{R}\times\mathbb{R}$.
@@ 4.1 Damped harmonic oscillator
Let us consider a one dimensional damped harmonic oscillator, such as in [2], whose continuous Lagrangian $L:{\mathsf{T}}\mathbb{R}\times\mathbb{R}\cong\mathbb{R}^{3}\to\mathbb{R}$ is given by
@@ 4.2 Ideal gas in a piston
Consider a perfect gas confined by a piston and contained in a cylinder, as was done by Gay-Balmaz and Yoshimura in [15]. In suitable units and for a suitable number of particles in the gas, the Lagrangian that describes this system is given by
@@ 4.3 Van der Waals gas in a piston
A more realistic example would be to consider a Van der Waals gas formed by particles without internal degrees of freedom, contained in a cylinder similar to that in the previous example. Considering a suitable value for the mass of the cylinder, as well as for the number of particles in the gas, the Lagrangian in suitable units is given by [23]: